### v2

# Spectrum Sharing in Satellite-Terrestrial Integrated Networks: Frameworks, Approaches, and Opportunities

Abstract
With the construction of low-earth orbit (LEO) satellite constellations, ubiquitous connectivity has been achieved. However, to accommodate the increasing communication needs in non-terrestrial networks (NTNs), wireless users in remote areas may require access to more spectrum than is currently allocated. Terrestrial networks (TNs), such as cellular networks, are deployed in specific areas, but many underused licensed spectrum bands remain in remote areas. Therefore, bringing NTNs to a shared spectrum with TNs can improve network capacity under reasonable interference management. However, in satellite-terrestrial integrated networks (STINs), the comprehensive coverage of a satellite and the unbalanced communication resources of STINs make it challenging to manage mutual interference between NTN and TN effectively. This article presents the fundamentals and prospects of spectrum sharing (SS) in STINs by introducing four SS frameworks, their potential application scenarios, and technical challenges. Furthermore, advanced SS approaches related to interference management in STINs and performance metrics of SS in STINs are introduced. Moreover, a preliminary performance evaluation showcases the potential for sharing the spectrum between NTN and TN. Finally, future research opportunities for SS in STINs are discussed.
I Introduction
Due to the boosting need for broader and more reliable wireless network coverage and data transmission, non-terrestrial networks (NTNs) have emerged as a promising solution. In urban areas with heavy traffic or remote areas where terrestrial networks (TNs) face geographical deployment limitations, NTNs can provide seamless access and ubiquitous connectivity [1]. Spectrum integration in satellite-terrestrial integrated networks (STINs) is a necessary step to improve the utilization efficiency of scarce spectrum resources. However, the coexistence of NTNs and TNs may be challenging, especially with regard to spectrum utilization and interference management. When NTN nodes such as low-earth orbit (LEO) satellites are incorporated into existing TNs via certain frequency bands, interference patterns can become more complicated due to different spatial propagation characteristics of signals. For example, the use of the Ka band or the Ku band for satellites may cause severe co-channel interference in existing communication systems. Moreover, within a fixed spectrum allocation, it is possible that while certain portions of the spectrum in NTNs are fully occupied, other segments within TNs may remain unutilized. The imbalance will lead to under-utilization of available spectrum resources and a significant hindrance to substantial improvements in system performance.
In recent years, the rapid growth of NTNs has paved the way toward realizing STINs. STINs, a promising candidate for future sixth-generation (6G) wireless communications, incorporate LEO nodes into existing TN to take advantage of NTN and meet diverse communication requirements [2]. By using STINs, the coverage, flexibility, and spectrum efficiency of the wireless network can be significantly promoted, as terrestrial users can access various TN and NTN nodes from vast geographical areas. Inevitably, the spectrum resource will become scarcer in STIN as more NTN nodes are introduced. This encourages us to manage the spectrum when TN and NTN co-exist in the STIN [3]. Meanwhile, the licensed spectrum to the static TN nodes cannot always be utilized efficiently, as spatial spectrum holes cannot be eliminated within the geographic area of interest. Hence, mobile NTN users can admit access to users within the spectrum holes and promote spectrum efficiency [4]. In existing work, one of the promising technologies for TN and NTN coexistence is spectrum sharing (SS), which allows the NTN to share the spectrum pre-allocated to the TN. SS allows the NTN to work in the same frequency band when the TN users are active, as long as its interference to the TN can be tolerated [4, 5]. In this way, the STIN can provide high-quality and seamless wireless services at a slight cost of performance loss under the practical restrictions of limited spectrum resources.
With the distinct advantages above, STINs can overwhelm conventional TNs in a variety of practical fields. As the key technique in STINs, SS becomes the main bottleneck in network performance promotion and faces several challenges [6]. First, a universal SS architecture cannot always meet the diverse needs of the STIN. Therefore, the SS frameworks for all possible STIN transmission scenarios should be carefully designed, including the uplink and downlink of both TN and NTN. Second, since interference in STIN leads to system performance degradation, mitigating interference becomes the most critical issue. Generally, such mitigation can be implemented by isolating users from the interference source and carefully managing the TN and NTN spectrum [7]. Meanwhile, the TN and NTN spectrum should be allocated appropriately, and the spectrum access scheme should be carefully designed [8]. In addition, since STINs are hierarchical and heterogeneous, the proper selection of performance metrics for system evaluation is also essential to the system, depending on the design objectives.
To address these concerns, we first consider the isolation of STIN users. One of the well-known approaches for this is the utilization of a protection zone for TN, where NTN users within the protection zone are not allowed to access the spectrum shared by TN and NTN. However, the appropriate definition of protection zones still remains a challenging problem. Meanwhile, some studies have been carried out for spectrum allocation and management. Typical work involves spatial SS, dynamic SS, and game-theoretic SS approaches, which are potentially beneficial for STINs. In spatial SS, the objective is to find the geographic areas that are good for NTN access in STINs [9]. In dynamic SS, the objective usually becomes finding the temporal holes in the spectrum resource by treating the NTN as a secondary network [10]. Both spatial SS and dynamic SS can be considered as frequency reuse schemes in STINs. For spectrum allocation and sharing, game models are competitive candidates in finding the optimal scheme by modeling TN and NTN as game players [11]. All of the above topics can be further studied to meet the performance requirements of future STINs.
FIGURE 1: An overview of SS in STINs.
This article focuses on the SS frameworks for various STIN scenarios and the state-of-the-art approaches to SS in STINs. The main contributions are summarized as follows.
  • We provide detailed discussions of possible SS frameworks in STIN and summarize their geographic and transmission characteristics, including all uplink and downlink scenario combinations for TN and NTN.
  • We investigate several state-of-the-art approaches for possible SS in STINs to promote spectrum efficiency and system utilization from the spatial, temporal, and frequency domains. Detailed illustrations of these approaches are also provided.
  • We summarize the potential performance metrics for evaluating the efficiency of SS in STINs and validate the improvement via a practical case study. Finally, future research opportunities for SS in STINs are highlighted.
II Frameworks of SS in STINs
In this section, we present four frameworks of SS in STINs as illustrated in Fig. 2. By implementing SS for NTN and TN in STINs, the total bandwidth is divided into two segments: 1) shared spectrum for both TN and NTN, and 2) reserved spectrum exclusively for NTN, as shown at the top of Fig. 2.
FIGURE 2: Four frameworks of SS in STINs.
II-A NTN Downlink and TN Downlink SS
When a satellite transmits signals to an NTN user and a BS transmits signals to a TN user, they can access the shared spectrum, as shown in Fig. 2(1). For example, this framework corresponds to the NTN satellite n256 frequency band and new radio (NR) n65 and n66 frequency bands in the International Telecommunication Union Radiocommunication Sector (ITU-R) and the 3rd Generation Partnership Project (3GPP) [12]. It’s worth noting that SS methods not only apply to the current spectrum regulations but also facilitate future spectrum utilization and standards development. In this case, TN users can be regarded as primary users and NTN users as secondary users. Designing a three-dimensional (3D) protection zone for each TN user is essential. The protection zone can be either spherical or tapered. Satellites outside this protection zone may utilize the shared spectrum while those within continue to use the reserved spectrum. This is because sharing spectrum between TN and NTN will not only enhance the communication performance of TN users through the expanded bandwidth but also result in interference. The protection zone is mainly set to protect the quality of service (QoS) of TN users. Theoretically, NTN users can also be affected by interference leakage from BSs, considering their beam directionality on the antenna surface in the horizontal plane. However, near-vertical transmissions from satellites can hardly be influenced. Moreover, satellites can mitigate mutual interference by constraining spectrum access within a specified distance range.
II-B NTN Downlink and TN Uplink SS
Consider a scenario involving sharing a reverse link spectrum in which an NTN user receives signals from a satellite. As depicted in Fig. 2(2), both the satellite and the NTN user can access the shared spectrum through TN uplink communications. In this framework, satellite spectrum sensing can be performed to detect the strength of TN uplink signals to reduce mutual interference. If the uplink signal exhibits a low level of strength, the NTN downlink transmission can be operated to prevent significant mutual interference. To be more specific, due to the directionality of satellite downlink signals, the corresponding interference to TN uplink transmissions on the near-horizontal plane is relatively small. In addition, owing to the extended propagation distance and limited antenna gain associated with the satellite’s side lobes, TN uplink users located within the reception range of these side lobes exert minimal interference on the satellites’ spectrum sensing capabilities. In particular, spectrum sensing is conducted because satellites and NTN users can neither determine TN users transmitting in the current time slot nor predict aggregated interference.
II-C NTN Uplink and TN Downlink SS
When an NTN user transmits signals to a satellite, they can opportunistically access the shared spectrum with TN downlink communications, as shown in Fig. 2(3). For example, this framework corresponds to the NTN satellite n256 frequency band and NR n2, n25, and n70 frequency bands in ITU-R. This framework is feasible due to a couple of factors. First, the BS is tilted downward for TN downlink communications and employs beamforming techniques to serve TN users. Because of the BS’s antenna angle and beamforming, interference leakage from the BS to the satellite is minimal. Second, during NTN uplink communications, the NTN user and the satellite can collaboratively perform spectrum sensing. This allows them to opportunistically access the spectrum while minimizing interference to and from TN users. It is important to note that, based on spectrum sensing conducted by NTN users, the aggregated interference power generated by NTN users is relatively lower than the desired signal power of the TN downlink communication. This is due to the high transmit power associated with TN BS, which also employs beamforming techniques.
II-D NTN Uplink and TN Uplink SS
In the ITU-R, NTN satellite n256 frequency band and NR n65 frequency band overlap. As shown in Fig. 2(4), allowing NTN uplink communications to share the spectrum with TN uplink communications is facing many problems, which makes it necessary for NTN to use the n256 band in areas where TN does not use the n65 band. First, numerous TN uplink users operate within a satellite’s broad coverage area, and the distances between these users and NTN uplink users can be significant. This results in limited detection capabilities, making it difficult for either type of user to perceive the other effectively. Second, TN users typically have a limited number of antennas, leading to considerable interference leakage from TN users in the vertical direction. This interference can overwhelm the desired signals from NTN uplink communications. Third, NTN users are likely to experience a high rate of missed detections regarding the presence of TN users. Consequently, this inability to accurately sense TN users can lead to significant interference from NTN uplink transmissions to TN communications. As a result, TN users’ communication performance often falls short of expectations. Sharing the spectrum between these two co-directional uplink links is unfavorable from the perspective of interference and sensing capabilities in many situations. However, this spectrum-sharing framework may be viable with appropriate spectrum-aware strategies when the number of TN uplink users is relatively small.
III Advanced Approaches to SS in STINs
This section presents various approaches for achieving SS in STINs under different scenarios.
FIGURE 3: Illustrations of advanced approaches to SS in STINs.
III-A Protection Zone-Based Approach
Method: Protection zones act as an auxiliary approach in SS as they define some geographical area in which users cannot access the spectrum of the primary system. SS techniques enable STIN users to sense and access the spectrum occupied by primary users to improve the spectrum efficiency. However, primary system interference cannot be ignored when users are close to interference sources. As protection zones are usually located around primary users, they can isolate interference for the users inside them, as shown in Fig. 3(a). While the users are outside the protection zones, they are free to access the spectrum.
Applications: With protection zones and knowledge of the user location in STINs, the SS technique enables intelligent monitoring of spatial spectrum holes and shares the spectrum outside the protection zone.
Difficulties and Solutions: Since the NTN is mobile and provides flexible spectrum access for users, finding the boundary for the protection zone becomes a challenging issue. The dynamics of STIN architectures also have significant impacts on the protection zones. Both challenges require statistical modeling and analysis based on user distributions.
III-B Spatial Spectrum Sensing (SSS)-Based Approach
Method: SSS helps secondary users access the spectrum that relies on sensing rather than knowing the positions of interfering nodes [13]. The term “spatial” refers to the ability of this method to capture the spatial characteristics of large-scale networks. The process begins with evaluating the interference power for a secondary user seeking to access a spectrum occupied by primary users. As shown in Fig. 3(b), if the test statistic for the interference power exceeds a predefined energy detection threshold, it indicates that the channel is currently busy. In this case, the secondary user transmits with a low probability. Conversely, if the interference power is below the threshold, the secondary user transmits with a high probability.
Applications: When SSS is implemented in STINs, NTN-related devices act as secondary users to monitor the busy status of the shared channel. The SSS-based approach suits the NTN downlink and TN uplink SS framework. Acquiring the locations of currently transmitting TN users at NTN satellites is difficult in practical engineering. By selecting an appropriate energy detection threshold, we can manage the interference generated by NTN toward TN.
Difficulties and Solutions: The primary technical challenge of this method is accurately determining the probabilities of false alarms and missed detections. It is crucial to control these probabilities following the Neyman-Pearson criterion. Possible solutions include calculating probability distributions based on the statistical properties of the interference under various events or using machine learning techniques to fit these probabilities.
III-C Joint Spatial Spectrum Sensing (JSSS)-Based Approach
Method: JSSS involves two different types of network nodes working together to perform SSS. As shown in Fig. 3(c), if both nodes determine that the channel is available for transmission, the secondary user will likely access the channel with a high probability. Conversely, if they disagree on availability, the secondary user will have a lower probability of accessing the channel.
Applications: This approach is particularly suitable for the NTN uplink and TN downlink SS framework. The reason is that the satellite, acting as the NTN receiver, may sometimes experience significant interference from the TN when the number of ground base stations is high in specific areas. Thus, the vertical interference leakage could not be ignored. As a result, the SSS performed by the satellite protects its communication from the interference generated by the TN. At the same time, the SSS done by the NTN user helps manage the interference produced by the NTN towards the TN.
Difficulties and Solutions: Since there are two different types of network nodes in JSSS, collaborating to decide whether SS can be accomplished has become a significant challenge. The energy detection thresholds of the satellite and the NTN user can be jointly optimized under the Neyman-Pearson criterion.
III-D Dynamic Spectrum Access-Based Approach
Method: Dynamic spectrum access (DSA) is a widely adopted SS approach that takes advantage of the temporal holes of the licensed spectrum. In STINs, users can sense and access the available spectrum as long as they are not causing significant interference to other users or the spectrum is idle, as shown in Fig. 3(d), and DSA can help achieve this objective by allowing secondary user access using the spectrum holes. In this way, the spectrum efficiency and system rate can thus be significantly promoted.
Applications: Since NTNs have high mobility, STIN users can adopt DSA strategies to access the NTN spectrum by integrating artificial intelligence (AI) tools. Similarly, NTNs can sense the spectrum licensed to TN users and adopt DSA to provide services for NTN users. Hence, the DSA protocol is essential for the DSA-based approach in SS and has been designed to guarantee the performance of STINs in existing studies.
Difficulties and Solutions: Since the NTN nodes need to sense and detect spectrum holes at various geolocations over time to provide wireless access, machine learning techniques incorporating dedicated geometric models are necessary for DSA in STINs. Meanwhile, since NTN nodes are independent, distributed NTN sensing schemes should be designed as well.
III-E Game Theory-Based Approach
Method: Game theory is a powerful theoretical approach that models the interacting users as rational players aiming to maximize their own benefits by selling and buying spectrum resources, as shown in Fig. 3(e). In STINs, users tend to make their decisions and compete with each other to access the shared spectrum, and ultimately decide the optimal SS pattern by obtaining the game Nash equilibrium.
Applications: In STINs, the spectrum allocation and access scheme of the TN/NTN nodes and the users can be described by adopting precise game models, and the optimal spectrum utilization pattern can be obtained. This is particularly suitable for NTN downlink scenarios, as potential users may access the shared spectrum when deciding their access mode.
Difficulties and Solutions: Since STINs are heterogeneous with multiple hierarchies, the internal coordination of various users and network devices in SS often exhibits considerably high complexity. In this sense, the selection of appropriate game theory approach is the key issue in modeling the sophisticated interaction behaviors of STINs.
IV Performance Metrics of SS in STINs
In this section, we introduce key metrics for evaluating the performance of SS in STINs from different perspectives.
IV-A User Throughput and Area Spectrum Efficiency in STINs
User throughput refers to the amount of data that the user can transmit and receive over a certain period, and system capacity is the sum of all user throughput. They both demonstrate the efficiency of STIN in continuously providing wireless service. Area spectrum efficiency (ASE) refers to the spectrum efficiency of a unit area, which can be used to evaluate the average performance of a large-scale network from a system-level perspective. Since STINs have dynamic architectures, a high and stable expected user throughput and ASE imply that STINs can provide high-quality wireless services, especially under large bandwidth and the assistance of SS.
IV-B Interference Intensity between NTN and TN
Interference intensity refers to the power strength per unit of bandwidth caused by secondary users to primary users, which denotes the negative impact of secondary users when accessing the shared spectrum in STINs. SS brings additional interference into the system, as users receive interference from the TN nodes during NTN downlink transmission, and the TN nodes receive interference from the users during NTN uplink transmission. Hence, a smaller interference intensity implies a higher efficiency and capacity of the STINs. Since the total interference power grows as bandwidth increases, the STIN performance can be guaranteed by maintaining a satisfactory interference intensity.
IV-C Spectrum Access Probability of NTN User
Spectrum access probability (SAP) refers to the likelihood that a secondary user will access the shared spectrum when they want to transmit packets. This metric indicates how fairly the spectrum is utilized among secondary users. When a channel is detected as busy, the SAP is low, prompting secondary users to either attempt to re-access the spectrum later or to utilize their reserved spectrum. For secondary users, such as NTN users or satellites within STINs, we can analyze or simulate the SAP when they employ an SS approach. Generally, a higher bandwidth is beneficial for SAP, as NTN users have more spectrum resources for SS.
IV-D End-to-End Latency in STINs
End-to-end (E2E) latency refers to the time a user receives a data packet across the network once the packet is triggered. This metric is essential in the SS of STINs, as it reflects the timeliness and fairness of packet transmission. In STINs, secondary users may need to delay their transmissions if they cannot access the spectrum in the current time slot. Additionally, when dealing with satellite-to-ground communication links, it is essential to consider propagation delay, which can add several milliseconds to the E2E latency. Generally, a larger bandwidth indicates a higher data rate and eventually results in a smaller E2E latency.
IV-E Energy Efficiency of NTN and/or TN
Energy efficiency (EE) refers to the number of bits a user can transmit by consuming per unit of power. As a key performance metric, EE is obtained by dividing data throughput by the amount of power consumption, which denotes energy utilization in transmitting signals. Since NTN nodes and users in STINs often face limited energy supply restrictions, it is crucial to promote their EE to transmit more data and provide better QoS. Meanwhile, increasing EE also enables us to better utilize scarce spectrum resources. Thus, a good guarantee of EE is essential for the overall system performance of STINs, and SS technique can naturally promote EE especially under large bandwidth conditions.
V Case Study: SS in STINs Using Protection Zone-based Approach
FIGURE 4: Simulation scenario.
FIGURE 5: System capacity and average data rate of STINs with SS and without SS. The total bandwidth is $300$ MHz, while the NTN bandwidth and the TN bandwidth are fixed to $120$ MHz and $180$ MHz, respectively, for STINs without SS. The BSs are randomly distributed in a $200$ km×$\times 200$ km remote area with a density of $\lambda_{\text{T}}=3\times 10^{-3}$/km2. Each BS has a service radius of $R_{\text{T}}=10$ km and transmit power of $P_{\text{T}}=46$ dBm. The pathloss exponent in TNs is $\alpha_{\text{T}}=3.5$. A typical TN user is assumed to be located at the center of the serving area, while typical NTN users are $50$ km away from the typical TN user. The noise power at the user side is $-110$ dBm. In NTNs, the satellites are distributed on an orbital sphere with a density of $\lambda_{\text{N}}=1\times 10^{-5}$/km2 at an altitude of $500$ km. The transmit power, main-lobe gain, side-lobe gain, and path loss exponent are $P_{\text{N}}=50$ dBm, $G_{\text{ml}}=30$ dBi, $G_{\text{sl}}=10$ dBi, and $\alpha_{\text{N}}\approx 2.0$, respectively.
In this section, we quantitatively study the effectiveness of SS in STINs by adopting the protection zone-based approach. The simulation scenarios are first depicted in Fig. 4. The NTN downlink and TN downlink framework is considered, where the NTN users are the primary users and the TN users are the secondary users. A protection zone-based approach is adopted, so that the BSs are located at least $R_{\text{p}}$ away from NTN users. For the case of SS in Fig. 4(a), NTN users are primary users each having a protection zone with radius $R_{\text{p}}$. An NTN user will use the reserved spectrum if there is a BS inside the protection zone and will use the shared spectrum if there is no BS inside. If SS is not adopted in STINs as shown in Fig. 4(b), NTN users and TN users will experience interference from the NTN spectrum and TN spectrum, respectively. This case study evaluates the aforementioned networks using two key performance metrics. First, the average data rate is determined by multiplying the product of spectral efficiency and bandwidth. Second, the system capacity of the NTN (or TN) is assessed by multiplying the spectral efficiency by the satellite (or BS) density, the surface area of satellites (or BSs), and the available bandwidth.
The simulation results are shown in Fig. 5, comparing the following three scenarios, i.e., Scenario 1: SS with the protection zone-based approach; Scenario 2: SS without the protection zone-based approach; Scenario 3: the setting without SS. Note that as reserved NTN bandwidth is considered in Scenario 2 and Scenario 3, the system capacity and average data rate of them remain almost the same. Specifically, in Fig. 5(a), the TN system capacity in Scenario 1 first surpasses that of Scenario 2 when $B_{\text{R}}$ is around $100$ MHz, but the latter subsequently exceeds it. After applying SS in Scenario 2, the excessive bandwidth can be utilized more efficiently, leading to an increase in sum and NTN system capacity compared to Scenario 3. The protection zone-based approach in Scenario 1 can utilize excessive reserved bandwidth and decrease co-frequency interference by setting an empirical protection radius $R_{\text{p}}$. If a BS is located within distance $R_{\text{p}}$, the NTN user will switch to the shared spectrum instead of the reserved spectrum. A gradual increase in both the sum and NTN system capacity can be observed.
In Fig. 5(b), the average data rate of the users in STINs with SS is compared to that of the users in STINs without SS. Compared to Scenario 3, there will be higher data rate for NTN user but lower data rate to TN user in Scenario 2, which is mainly due to stronger co-frequency interference from satellite networks. In Scenario 1, with an increase of $B_{\text{R}}$, the data rate of the NTN user and the TN user using shared spectrum decreases because more bandwidth is allocated to the reserved spectrum. However, the data rate of the NTN user using reserved spectrum increases. Therefore, by selecting the appropriate reserved bandwidth according to implementation requirements, a win-win solution can be obtained by using SS with the protection zone design.
VI Future Opportunities of SS in STINs
In this section, we briefly discuss future applications of SS in STINs. An overview of these potential applications is presented in Fig. 6, and detailed descriptions of each of them are given as follows.
FIGURE 6: Future opportunities of SS in STINs.
VI-A SS for Multi-Tier NTN and TN
Due to the limited availability of orbital positions, satellites are deployed at various altitudes. Different types of satellites have varying antenna patterns and transmit powers, and the number of satellites in different tiers also varies. As a result, the single-tier STIN will evolve into a multi-tier heterogeneous STIN (HetSTIN). However, this multi-tier HetSTIN faces a significant challenge: the conflict between limited spectrum resources and the increasing demand for satellite services. Effectively sharing the spectrum among the multi-tier NTN and multi-tier TN is crucial for successfully integrating satellite and terrestrial networks. Addressing the complexities of interference management and the imbalanced communication resources that arise from the heterogeneous nature of multi-tier HetSTINs is essential in the design of SS.
VI-B SS for Multi-Connectivity-Based Cooperative STINs
Multi-connectivity in STINs, which has been referred to in the 3GPP Technical Standard [14], indicates that a user can simultaneously connect to multiple network nodes, including BSs or access points in TNs and satellites in NTNs. With proper synchronization and mitigation of delay, phase misalignment, and Doppler shifts with algorithms [15], cooperative communication, e.g., relaying, joint transmission, or reception, can be utilized to connect different network tiers in STINs and promote more resilient and adaptable communications. However, the bandwidth allocated to a user, which may have multiple connection links, can be restricted due to the scarcity of available spectrum resources. Hence, the allocation of spectrum resources among multiple users, each possessing multiple links, presents significant research opportunities. In addition, the organization of cooperative network nodes, whether centralized or decentralized, requires further exploration and investigation.
VI-C Multiple Access-Enabled Cognitive SS in STINs
Multiple access allows multiple users to be served simultaneously in the same frequency band. Using this characteristic, multiple access and cognitive SS can be connected by sensing the real-time radio environment. The cognitive radio can sense the surroundings and adjust the transmission parameters accordingly. With these sensing results, secondary users can assess the unutilized spectrum without interfering with primary users. In addition, this fusion also offers some other benefits, including increased network capacity, more efficient and flexible resource allocation, and decreased interference. By continuously monitoring the radio environment and adjusting transmission parameters, more reliable network performance can be ensured, and a more equitable distribution of spectrum resources among users can be promoted.
VI-D SS for Different Satellite Constellations
As satellite constellations have expanded to mega size in recent years, the number of different satellite constellations built by different satellite companies is also increasing. Satellite constellations can also share their proprietary spectrum resources to improve the system’s capacity. Due to the mobile nature of the satellites, the NTN architecture is always dynamic. It changes rapidly, leading to the significant challenge of maintaining reliable and high-quality services under the restriction of limited and congested spectrum resources. Different satellite constellations face their transmission conditions as service providers. They require compatible spectrum access strategies. SS techniques enable the high-efficiency utilization of spectrum resources while introducing controllable interference. Meanwhile, to meet the constellation dynamics, the designed SS strategy should also be adaptive to available satellites and spectrum resources from time to time in providing wireless services.
VII Conclusions
This article investigated the frameworks, approaches, performance metrics, and future research opportunities for SS in STINs. We introduced four general SS frameworks in STINs, including different uplink and downlink combinations for NTN and TN, and discussed their characteristics. Then, state-of-the-art approaches for SS to reuse spectrum resources in space, time, and frequency domains in STINs were comprehensively reviewed. More importantly, related metrics were presented to study the SS techniques further and evaluate the performance of the system and the individual user. A case study on the protection zone-based approach for SS in STINs was carried out, and the simulation results showed that a win-win solution could be obtained for improved system capacity. Finally, we provided promising opportunities and insightful future research on SS in STINs.
Biographies
Bodong Shang (bdshang@eitech.edu.cn) received his M.S. degree from Xidian University, China, and his Ph.D. degree from the Department of Electrical and Computer Engineering at Virginia Tech, Blacksburg, USA, in 2021, and he was a Postdoctoral Research Associate at Carnegie Mellon University, Pittsburgh, USA. He held a research internship position at Nokia Bell Labs, USA. He is an Assistant Professor at Eastern Institute for Advanced Study, Eastern Institute of Technology (EIT), Ningbo, China. His research interests include space-air-ground-sea integrated networks, non-terrestrial networks, and space information networks.
Zheng Wang (zwang@idt.eitech.edu.cn) received his Ph.D. degree from the Department of Electrical and Computer Engineering at George Mason University, Fairfax, USA, in 2020, and he was a Postdoctoral Research Associate at Virginia Tech, Blacksburg, USA. He is a Research Associate Professor at Ningbo Institute of Digital Twin, Eastern Institute of Technology (EIT), Ningbo, China. His research areas include satellite-terrestrial integrated networks, game theory, and physical layer communication.
Xiangyu Li (xyli@eitech.edu.cn) received an M.S. degree in Electrical and Computer Engineering from Georgia Institute of Technology, Atlanta, USA, in 2023, and he is currently pursuing a Ph.D. degree with Shanghai Jiao Tong University (SJTU), Shanghai, China, at the Eastern Institute of Technology (EIT), Ningbo and SJTU Joint PhD Program. His research interests include space-air-ground integrated networks, non-terrestrial networks, and performance analysis of wireless systems.
Chungang Yang (guideyang2050@163.com) is a full professor at Xidian University, leading the GUIDE: Game, Utility, Artificial Intelligent Design for Emerging Communications research team. His research interests include artificial intelligence 6G wireless mobile networks, intent-driven networks, space-terrestrial networks, and game theory for emerging communication networks.
Chao Ren (chaoren@ustb.edu) is an Associate Professor and Deputy Director with the University of Science and Technology in Beijing, China. He is an Editorial Board Member for the Journal of Communications and Frontiers in Computer Science, recognized with an Outstanding Editor Award. He has been a Guest Editor for Space-Integrated-Ground Information Networks and Information Security Research, Co-Chair of the UCOM 2023-2025 Workshop, and Technical Program Committee member for over 30 international conferences in communications.
Haijun Zhang (haijunzhang@ieee.org) is a full professor and dean at the University of Science and Technology in Beijing, China. He serves/served as an Editor of IEEE Transactions on Information Forensics and Security, IEEE Transactions on Communications, IEEE Transactions on Network Science and Engineering, and IEEE Transactions on Vehicular Technology. He received the IEEE CSIM Technical Committee Best Journal Paper Award in 2018, the IEEE ComSoc Young Author Best Paper Award in 2017, and the IEEE ComSoc Asia-Pacific Best Young Researcher Award in 2019. He is a Fellow of IEEE.